OpenStudy (anonymous):

Which of the following equations is the result of completing the square on x^2 - 6x - 9 = 0?

OpenStudy (anonymous):

(x - 3)2= 0 (x - 3)2= 9 (x - 3)2= 18

OpenStudy (anonymous):

(x - 3)^2= 0 (x - 3)^2= 9 (x - 3)^2= 18

OpenStudy (anonymous):

Third one.